OpenStudy (anonymous):

Simplify |10-(-5)| my wife and I are helping our daughter with pre algebra and we both came up with diferrent answer to the equation. I say the answer is 5 and she says it is 15.

OpenStudy (anonymous):

15

OpenStudy (anonymous):

You multiply the negative first, so it would be 15.

OpenStudy (anonymous):

\[a-(-b)=a+b\]

OpenStudy (precal):

10-(-5) =15 the absolute value means what is the distance from 0 to 15. The answer is 15 units.